### v1

# TiRE-GAN: Task-Incentivized Generative Learning Models for Radiomap Estimation with Radio Propagation Model

Abstract
Enriching geometric information on radio frequency (RF) signal power distribution in wireless communication systems, the radiomap has become an essential tool for resource allocation and network management. Usually, a dense radiomap is reconstructed from sparse observations collected by deployed sensors or mobile devices, which makes the radiomap estimation an urgent challenge. To leverage both physical principles of radio propagation models and data statistics from sparse observations, this work introduces a novel task-incentivized generative learning model, namely TiRE-GAN, for radiomap estimation. Specifically, we first introduce a radio depth map as input to capture the overall pattern of radio propagation and shadowing effects, following which a task-driven incentive network is proposed to provide feedback for radiomap compensation depending on downstream tasks. Our experimental results demonstrate the power of the radio depth map to capture radio propagation information, together with the efficiency of the proposed TiRE-GAN for radiomap estimation.
Index Terms:
radiomap estimation, generative adversarial networks, radio propagation model
publicationid: pubid: 978-1-6654-3540-6/22 © 2022 IEEE
I Introduction
The expansion of sensor networks and edge computing systems has stimulated many new technologies in the next generation of wireless communications, such as vehicle-to-everything (v2x) communications [1], unmanned aerial vehicle (UAV)-assisted networking [2] and fifth-generation (5G) terrestrial and non-terrestrial network (NTN) services [3]. All these applications require an accurate description of radio frequency (RF) spectrum coverage and efficient assessment of the radio environment, which lead to the new concept of radiomaps [4]. Describing the RF power spectral density (PSD) over different geometric locations, frequencies, and time, radiomaps can enrich the information of spectrum coverage and assist the aforementioned spectrum management applications [5]. Usually, dense radiomaps are reconstructed from sparse observations collected by mobile sensors and user devices, as shown in Fig. 1. To fully leverage the power of radiomaps in spectrum resource allocation and network optimization, efficient radiomap estimation (RME) from sparse observations has emerged as an urgent challenge.
Existing methods of radiomap estimation can be categorized into model-based or learning-based approaches, each bearing its distinct constraints. Model-based methods usually assume a specific radio propagation model, with which unknown parameters are estimated by minimizing the error between observed samples and predictions. Typical examples include the log-distance path loss (LDPL) models [6], inverse distance weighted interpolation [7] and thin-plate splines kernels [8]. However, the performances of model-based RME highly rely on the selection of physical radio propagation models, which are usually unknown in practical applications. Moreover, the PSD distribution sometimes is more sensitive to surrounding obstacles rather than following a strict empirical model, especially in complicated environments. The efficient extraction of radio model information from realistic data samples remains an open question.
With the development of artificial intelligence (AI), recent interests of RME focus on learning-based methods [9]. In [10], a fast radiomap estimation is proposed based on Unet. Another work utilizing deep convolutional neural networks is [11], where an auto-encoder structure is presented for radiomap estimation. Other learning-based methods also include the conditional generative adversarial nets (cGAN) [3] and transformers [12]. Despite many successes, efficient learning-based RME requires sufficient data samples with satisfying quality, which are usually inaccessible in realistic scenarios. Although some works [13, 14] consider integrating the radio propagation models with learning machines to address the data scarcity, the efficient model selection remains challenging. Moreover, how to satisfy the special requirements of downstream tasks, such as outage detection and network planning, is another critical concern.
To address the aforementioned problems, this work proposed a Task-incentivized Radiomap Estimation based on Generative Adversarial Networks (TiRE-GAN) for radiomap estimation, leveraging the synergies of radio propagation model and generative AI. Specifically, a radio depth map is introduced as an additional input channel of the generator to capture the overall radio propagation information and shadowing effects. Then, a task-incentivized block is proposed to provide feedback for radiomap generation to compensate for the estimated PSD based on downstream tasks. Our contributions can be summarized as follows:
  • We introduce a novel radio depth map (RDM) as input conditions to the cGAN, serving as a proxy for the shadowing effects and radio propagation behavior, thereby infusing the learning process with additional insights into the physical principles of radio propagation.
  • We integrate the physics-conditioned cGAN with a novel task-incentivized network to provide compensation to the estimated radiomaps depending on downstream tasks, which significantly enhances the RME performance.
  • Our experimental results demonstrate the power of the proposed radio depth map in characterizing radio propagation behavior, as well as the efficiency of the proposed TiRE-GAN in radiomap estimation.
II Problem Description
Figure 1: Problem Description of RME: Given the distribution of buildings, transmitter locations (in read), and the sparse observations collected by mobile sensors and devices, we aim to predict the dense radiomaps.
Figure 2: Overall Structure of TiRE-GAN consisting of 3 parts: 1) The model-based data preprocessing module utilizes the input observed information (e.i. building map and transmitter location) to generate the radio depth map and then stacks the four layers together as input; 2) The radiomap (RM) generation module is in a cGAN generator structure, which learn and produce the prediction of radiomap; 3) The feedback module consists of 2 parts: a cGAN discriminator and a task-incentivized network. Both modules produce a loss term as feedback to improve the generator.
In this section, we first introduce our problem formulation. As shown in Fig. 1, the radiomap estimation is formulated as a sparse sensing problem, where a dense radiomap $\hat{M}_{R}$ is reconstructed given the landscape information of the surrounding environment, sparse observations collected from sensors and mobile devices, and locations of transmitters, formulated by the following equation:
$\hat{M}_{R}=g(M_{U},M_{T},M_{S},\mathcal{X}),$ (1)
where $M_{U}$, $M_{T}$, $M_{S}$ and $\mathcal{X}$ represent the landscape map, the encoded transmitter location, the sampled radiomap, and additional extracted features, respectively.
Different from the interpolation formulation in [6] and [7], where the unobserved PSD values are estimated only using the observed samples in the same radiomap, we follow the classic prediction formulation and assume a set of training pairs of $(M_{R},\mathcal{F})$ with =$\mathcal{F}=\{M_{U},M_{T},M_{S},\mathcal{X}\}$ to train the mapping function $g(\cdot)$ by
$\min_{g(\cdot)}\quad L(M_{R},\hat{M}_{R}),$ (2)
where $L(\cdot)$ refers to the designed loss function.
In this work, our objectives can be summarized as two parts: 1) design an effective mapping function $g(\cdot)$ by integrating the cGAN with a novel task-incentivized network; and 2) design a suitable loss function $L(\cdot)$ to embed the physical principles of radio propagation. We shall introduce the details of our proposed TiRE-GAN in Section III.
III Methods
III-A Overall Structure of TiRE-GAN Framework
We now introduce the overall structure of our proposed TiRE-GAN. As shown in Fig. 2, the architecture of the proposed framework can be structured into three components: 1) model-based data preprocessing; 2) radiomap generation based on cGAN; and 3) task-incentivized feedback.
The overall operational sequence of the training process is delineated as follows. To embed the radio propagation model, the urban landscape map and the transmitter location data are processed to fabricate a physics-embedded Radio Depth Map (RDM), which will be further discussed in Section III-B2. Subsequently, this RDM, alongside the sampled radiomap, urban map, and transmitter location data, are amalgamated into a quadruple-channel input for the cGAN. Throughout the training phase, a pre-trained task-incentivized network is incorporated as a contributory factor within the loss function for the cGAN’s generator, enhancing the model’s ability to compensate for the radiomap estimations by embedding physics-based constraints and the rewards from downstream tasks. We will then introduce each module as follows.
III-B Model-Based Data Preprocessing
The data preprocessing module consists of two parts: 1) process the observed features $\{M_{U},M_{T},M_{S}\}$ into aligned inputs of cGAN; and 2) generate a radio depth map embedding the radio propagation behavior.
III-B1 Observed Data Preprocessing
Suppose that the radiomap is formatted in a $N\times N$ grid, i.e., $M_{R}\in\mathbb{R}^{N\times N}$. The landscape map of buildings is encoded as a binary segmented image $M_{U}\in\mathbb{R}^{N\times N}$, where building areas are marked by 1 and open areas are annotated by 0. Similar to the building map, we encode the transmitter locations into a binary image $M_{T}\in\mathbb{R}^{N\times N}$, where the transmitter locations are marked as 1; otherwise, the pixel values are zero. The sparse observations can be viewed as the corresponding pixels in gray-scale images $M_{S}\in\mathbb{R}^{N\times N}$. For convenience, we first use a threshold to eliminate the extremely low-value observations, and then normalize the PSD value into $0\sim 1$ for the observed locations. Since linear normalization is applied, we can easily re-scale the estimated radiomap back to the original data domain. For the missing values in $M_{S}$, we set them as zeros. Since a convolutional neural network is applied for the cGAN structure, zero-padding does not impact the final radiomap estimation.
III-B2 Radio Depth Map
To embed the radio propagation information, we introduce the Radio Depth Map (RDM), $\mathcal{X}=M_{D}\in\mathbb{R}^{N\times N}$ as a novel model-based attribute, derived from the distribution of buildings and the spatial locations of transmitters. Similar to a depth map in computer vision which provides information on the distances of surfaces from a specific viewpoint, the RDM is formulated as:
$M_{D}(x,y)=n\left(\sum_{t=1}^{N_{T}}P_{t}(x,y)\cdot B_{t}(x,y)\right),$ (3)
where $(x,y)$ denotes the two-dimensional coordinates of the target location, $P_{t}(x,y)$ denotes the distance-based path loss gain from the $t$-th transmitter, and the $B_{t}(x,y)$ capture the impact of buildings from the viewpoint of $t$-th transmitter.
Suppose that $C(l,l_{t})$ identifies the set of pixels in the direct line of sight between the target location $l=(x,y)$ and location $l_{t}=(x_{t},y_{t})$ of the $t$-th transmitter. The shadowing term impacted by buildings is calculated by
$B_{t}(x,y)=\frac{\sum_{i\in C(l,l_{t})}(1-M_{U}(i))}{\sum_{i\in C(l,l_{t})}1},$ (4)
which characterizes the ratio of non-building pixels between $l$ and $l_{t}$. $B_{t}(x,y)$ is larger and the shadowing effect is smaller if the radio propagates through fewer buildings.
To capture the path loss $P_{t}(x,y)$, we employ the Inverse Distance Weighting (IDW) model, calculated by
$P_{t}(x,y)=d(l,l_{t})^{-\lambda},\ \lambda>0,$ (5)
where $d(l,l_{t})$ denotes the Euclidean distance between the target point $l$ and the $t$-th the transmitter. The parameter $\lambda$ governs the rate, at which signal strength diminishes with increasing distance, necessitating a positive value to ensure physical plausibility in the depth map. An exemplary radio depth map is shown in the orange block of Fig. 2.
III-C cGAN Network
In this work, we employ a conditional Generative Adversarial Network (cGAN) [15] as the foundational architecture, which consists of two principal components: a generator $G$ and a discriminator $D$. Different from the traditional GAN which generates data from random noise, the generator in cGAN can generate data from given features $\mathcal{F}$ and incorporate additional information $\mathbf{I}$ in the discriminator. In our radiomap estimation, the conditional features consist of four channels as processed in the Data Preprocessing Module, i.e., =$\mathcal{F}=\{M_{U},M_{T},M_{S},M_{D}\}$ as shown in Fig. 2. The additional information $\mathbf{I}$ can be designed as a function of the region features, i.e., $\mathbf{I}=h(\mathcal{F})$ or some determined labels, such as fake or real [16]. Then, the basic cGAN can be formulated as
$\min_{G}\max_{D}\quad V(D,G)=\mathbb{E}_{y,I\sim p_{\rm data}(\bf{y},\bf{I})}[\log D(\bf{y,I})]+$
$\mathbb{E}_{f\sim p_{\rm data}(\bf{\mathcal{F}})}[\log(1-D(G(\bf{\mathcal{F}}),\mathbf{I}))],$ (6)
where $p_{\rm data}(\bf{y,I})$ is the joint distribution with the prior knowledge $\mathbf{I}$, and $\mathbf{y}$ is the generated data. Generally, the optimization of a min-max problem in cGAN can be split into two sub-problems, and train $D$ and $G$ iteratively. Specifically, the loss function to optimize $D$ is
$L_{D}=-V(D,G),$ (7)
and the general loss function to optimize $G$ is
$L_{G}=\mathbb{E}_{x\sim p_{\rm data}(\bf{\mathcal{F}})}[\log(1-D(G(\bf{\mathcal{F}}),\mathbf{I}))].$ (8)
To leverage both radio propagation model information and data statistics from observed samples, in this work, we will customize the loss function of the generator as
$L=L_{G}+\alpha L_{MSE}+\beta L_{R}.$ (9)
Here, $L_{MSE}$ describes the error of RME by Mean Squared Error (MSE), i.e.,
$L_{MSE}(\mathbf{y},\hat{\mathbf{y}})=\frac{1}{K}\sum_{i=1}^{K}(y_{i}-\hat{y}_{i})^{2},$ (10)
where $K$ is the total number of pixels, $\mathbf{y}$ and $\hat{\mathbf{y}}$ are the ground truth and the predicted radiomap, while $y_{i}$ and $\hat{y}_{i}$ are the true and predicted pixel value of the $i$th pixel. $L_{R}$ is a task-incentivized regularization term from the feedback module, which will be introduced in Section III-D.
III-D Feedback Design
The feedback module consists of two components: 1) a traditional cGAN discriminator; and 2) a task-incentivized network (TIN) based on downstream tasks.
III-D1 cGAN Discriminator
Conversely to the cGAN generator, the discriminator aims to detect whether the generated data is fake or not. Here, we consider three-channels as the input of the discriminator, where two are expanded image-sized one-hot encodings representing the binary truthfulness of the input, and the third channel is the generated radiomap. The discriminator outputs a scalar value, indicating the likelihood of whether the provided radiomap is genuine. The discriminator is trained by the binary cross entropy (BCE) loss as
$\vspace{-5mm}L_{D}(\mathbf{y},\hat{\mathbf{y}})=-\frac{1}{K}\sum_{i=1}^{K}\left[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right].$ (11)
Figure 3: Task-incentivized network (TIN) structure: We design the TIN based on Unet, which contains 4 encoding blocks and 4 decoding blocks. Each encoding block consists of 3 convolution layers and a pooling layer. Each decoding block consists of one concatenation layer, a transpose convolution layer, and 3 convolution layers.
III-D2 Task-Incentivized Network
In traditional RME, the objective transcends the mere accuracy of radio signal strength at discrete locations within a dense radiomap. However, in practical scenarios, the final objective of RME is to assist spectrum planning and network management. For example, one can easily identify the outage of regions based on their radiomap by setting a certain threshold of service failure. To compensate for the radiomap estimations based on downstream tasks, we introduce a TIN module as shown in Fig. 2.
As an exemplary TIN, we adopt outage detection as a downstream task. More specifically, the proposed TIN utilizes a UNet architecture for semantic segmentation of radio power outage as Fig. 3, where the input constitutes a radiomap and the output is an outage map indicating signal coverage or absence. The training of the TIN is conducted using ground-truth radio coverage and outage maps, thereby ensuring that the network learns to accurately predict outage areas based on real-world data. Once trained, the TIN operates within the cGAN framework as a pre-trained component. During inference, the output generated by the cGAN’s generator serves as the input to the TIN, enabling it to contribute an additional term $L_{R}$ to the loss function of the cGAN. This term is specifically designed to encapsulate aspects of human-centric interest, such as the distribution and characteristics of signal outages, thereby guiding the cGAN toward generating radiomaps that are not only precise but also aligned with the real-world considerations and requirements of downstream tasks. This integration of the TIN within the cGAN architecture exemplifies a strategic convergence of machine learning precision with domain-specific insights, enhancing the utility and applicability of the generated radiomaps.
The TIN provides the task-incentivized regularization term $L_{R}$ in the loss function of the generator, which is calculated by the pre-trained TIN using the MSE loss to measure the error between the predicted and the ground truth outage maps:
$L_{R}(z,\hat{z})=\frac{1}{K}\sum_{i=1}^{K}(z_{i}-\hat{z}_{i})^{2},$ (12)
where the $z$ and $\hat{z}$ are the ground truth and the predicted outage map, while $z_{i}$ and $\hat{z}_{i}$ are the true and predicted pixel values of the $i$th pixel. $K$ is the number of pixels.
Note that, our TIN can be integrated with any RM-assisted applications. Some other typical examples also include ray tracing or wireless localization.
Figure 4: Performance comparison (sr=1) between Kriging, matrix completion(MC), AutoEncoder (AE), RadioUNet (RU), conditional GAN (cGAN), our TiRE-GAN, and the ground truth (GT) on 3 radiomap examples. TiRE-GAN achieves the best radio propagation and the shadowing effect prediction.
IV Experiments
IV-A Experiment Setup
IV-A1 Data description
In this work, we utilize the RadiomapSeer dataset [10] for evaluation, which comprises an extensive set of 700 maps from a variety of metropolitan areas, including Ankara, Berlin, Glasgow, Ljubljana, London, and Tel Aviv. Each map delineates a quadrant of $256\times 256$ square meters, depicted with a granularity of $256\times 256$ pixels. The edifices are encoded by assigning a pixel intensity of one, whereas open spaces are represented by a zero value. Furthermore, the precise locations of transmitters are encoded with a pixel intensity of one superimposed on a null background, facilitating accurate localization.
IV-A2 Method Comparison
To evaluate the performance of our TiRE-GAN, we compare with both state-of-the-art (SOTA) model-based methods, including Kriging [17], and data-driven approaches, including Matrix completion (MC) [18], AutoEncoder (AE) [11], RadioUnet (RU) [10] and conventional cGAN [15]. For the interpolation-based method, we reconstruct the radiomap with observations in the same radiomap. For the prediction-based method, we split the dataset as 4000/800/800 (70%/15%/15%) for training/validation/testing sets. In our TiRE-GAN, we set $\alpha=1$ and $\beta=0.1$ in Eq. (9).
IV-B Overall performance
We first present the overall performance of different radiomap estimations under different sampling ratios (sr). In this work, we evaluate the performance based on the normalized radiomaps via MSE denoted by
$\text{MSE}(\mathbf{y},\hat{\mathbf{y}})=\frac{1}{K}\sum_{i=1}^{K}(y_{i}-\hat{y}_{i})^{2},$ (13)
and normalized MSE (NMSE) calculated by
$\text{NMSE}(\mathbf{y},\hat{\mathbf{y}})=\frac{1}{K}\sum_{i=1}^{K}\frac{(y_{i}-\hat{y}_{i})^{2}}{y_{i}^{2}},$ (14)
where $K=256^{2}$. Note that, since the linear normalization is applied, we can easily re-scale the estimated radiomap back to the original data domain. The results are presented in Table I and II. From the results, our proposed TiRE-GAN always achieves superior performance against all other SOTA approaches. Specifically, with the assistance of our proposed radio depth map, our method can achieve a very high reconstruction accuracy even with a very small sampling ratio and insufficient training samples, which validates the effectiveness of the proposed method. The visualization results of different methods in 3 radiomap examples are shown in Fig. 4, where the sample ratio is 1%. From the visualization results, the radioUnet, cGAN, and TiRE-GAN have better performances than Kringing, MC, and AE in capturing the overall radio propagation patterns. Compared to radioUnet and cGAN, our TiRE-GAN has more explicit boundaries for the shadowing effects, resulting from the compensation from the TIN feedback. The numerical results, together with the visualization results demonstrate the efficiency of our proposed TiRE-GAN for radiomap estimation.
TABLE I: Test NMSE Comparison (×$\times 10^{-2}$)
sr (%) 0.03 0.1 0.5 1 3 5 10 Average
Krigin 24.5696 18.9280 14.2451 12.9608 14.7171 11.3540 10.9647 15.3913
RU 6.5421 5.1111 2.6374 2.2206 1.8192 1.5191 1.2704 2.8744
AE 54.7656 53.8440 45.6640 43.3302 26.0041 20.7167 14.3275 36.9503
AE deep 50.6597 48.8197 21.4272 16.2537 6.8252 6.5503 5.8179 156.3537
cGAN 11.5754 5.4805 2.1345 1.8301 1.3009 1.1894 1.0295 3.5057
MC 83.0969 74.7810 69.0133 66.9996 62.8215 56.3645 55.4314 66.9297
TiRE-GAN 1.9866 1.2608 0.7282 0.6084 0.5228 0.5119 0.4615 0.8686
TABLE II: Test MSE Comparison (×$\times 10^{-4}$)
sr (%) 0.03 0.1 0.5 1 3 5 10 Average
Krigin 188.7554 141.3767 110.7954 101.5476 91.9940 89.0443 86.0208 115.5340
RU 24.6458 19.3893 10.0221 8.4693 7.0856 5.9640 4.9978 11.5105
AE 230.6564 226.6141 194.4702 186.5724 113.0439 87.7205 65.0134 157.7272
AE deep 209.2621 201.4013 91.9589 70.0012 26.8177 26.4321 23.4719 92.7636
cGAN 45.8797 22.0908 8.4572 7.2467 5.0700 4.6442 4.0541 13.9203
MC 342.1892 307.9448 284.1937 275.9010 258.6960 232.1065 228.2640 275.6136
TiRE-GAN 7.0604 4.6453 2.7678 2.3472 2.0377 2.0015 1.8137 3.2390
IV-C Robustness Against Noise
To evaluate the robustness of TiRE-GAN, we conduct the experiment, where a noisy input pattern is substituted for the original sparse observation channel by introducing additive Gaussian noise to each sampled value in the radiomap. To showcase the effectiveness of the proposed radio depth map and TIN blocks, we compare our TiRE-GAN with Kriging, matrix completion(MC), AutoEncoder (AE), RadioUNet (RU), and conditional GAN (cGAN) under different signal-to-noise ratios (SNR) as Fig. 5. Here, we consider a sampling ratio of 1% for all the test cases. From the results, our TiRE-GAN always has better performance. Moreover, the MSE of TiRE-GAN increases slower than those of other methods, especially compared to the high performance methods (RadioUNet, cGAN), demonstrating the robustness of the TiRE-GAN against noise.
Figure 5: The MSE (×$\times 10^{-4}$) of Kriging, matrix completion(MC), AutoEncoder (AE), RadioUNet (RU), conditional GAN (cGAN), and TiRE-GAN of different SNRs under a sample ratio=1%.
IV-D Outage Evaluation of Generated radiomap
To further evaluate the performance of reconstructed radiomaps in downstream tasks, we compare the performance of RM-assisted outage detection. Here, we input the generated radiomaps from different methods into the pre-trained outage detection networks. The numerical results of the whole datasets with sr=1% and the exemplary visualization results are shown in Table III and Fig. 6, respectively. From the results, the traditional model-based methods and conventional auto-encoder perform poorly in the outage detection. Our TiRE-GAN always has superior performance, which demonstrates the effectiveness of TIN in radiomap compensation for downstream tasks.
TABLE III: Outage Map MSE comparison
Method Kriging MC AE
MSE 0.078134 0.075765 0.049377
Method RU cGAN TiRE-GAN
MSE 0.028439 0.030352 0.026682
Figure 6: Illustration of Outage Map Prediction under sr=1%. We compared the outage map predicted by the generated radiomap from Kriging, matrix completion(MC), AutoEncoder (AE), RadioUNet (RU), conditional GAN (cGAN), and our TiRE-GAN. The corresponding radiomap and the ground truth outage map(GT) are shown on the right for reference. The result shows that TiRE-GAN has the best capability to predict the outage map with a clear boundary and range.
V Conclusion
In this work, we introduce a task-incentivized generative learning model, namely TiRE-GAN for radiomap estimation. More specifically, to characterize the overall radio propagation behavior, we introduce a radio depth map as additional physics-embedded features as the condition to guide the data-driven generative models. To compensate for the radiomap estimation for downstream tasks, we introduce a task-incentivized network to provide feedback to predict the detailed features in the radiomap. From our experimental results, the proposed TiRE-GAN always achieves superior performance, even with insufficient training samples, which addresses the barriers of conventional learning approaches. Our numerical and visualization results demonstrate the effectiveness of the proposed radio depth map and TiRE-GAN. Future research endeavors will delve into the optimization of the generator’s reward mechanism. Another promising direction is the multi-band radiomap estimation for different frequencies.

### v3

@@
-TiRE-GAN: Task-Incentivized Generative Learning Models for Radiomap Estimation with Radio Propagation Model
+TiRE-GAN: Task-Incentivized Generative Learning for Radiomap Estimation
 Abstract
-Enriching geometric information on radio frequency (RF) signal power distribution in wireless communication systems, the radiomap has become an essential tool for resource allocation and network management.
-Usually, a dense radiomap is reconstructed from sparse observations collected by deployed sensors or mobile devices, which makes the radiomap estimation an urgent challenge.
+To characterize radio frequency (RF) signal power distribution in wireless communication systems, the radiomap is a useful tool for resource allocation and network management.
+Usually, a dense radiomap is reconstructed from sparse observations collected by deployed sensors or mobile devices.
 To leverage both physical principles of radio propagation models and data statistics from sparse observations, this work introduces a novel task-incentivized generative learning model, namely TiRE-GAN, for radiomap estimation.
-Specifically, we first introduce a radio depth map as input to capture the overall pattern of radio propagation and shadowing effects, following which a task-driven incentive network is proposed to provide feedback for radiomap compensation depending on downstream tasks.
-Our experimental results demonstrate the power of the radio depth map to capture radio propagation information, together with the efficiency of the proposed TiRE-GAN for radiomap estimation.
+Specifically, we first introduce a radio depth map to capture the overall pattern of radio propagation and shadowing effects, following which a task-driven incentive network is proposed to provide feedback for radiomap compensation depending on downstream tasks.
+Our experimental results demonstrate the power of the radio depth map to capture radio propagation information, and the efficiency of the proposed TiRE-GAN for radiomap estimation.
 Index Terms:
 Radiomap estimation, generative adversarial networks, radio propagation model
-publicationid: pubid: 978-1-6654-3540-6/22 © 2022 IEEE
+publicationid: pubid: $0$–$00$ 2021 IEEE
 I Introduction
-The expansion of sensor networks and edge computing systems has stimulated many new technologies in the next generation of wireless communications, such as vehicle-to-everything (v2x) communications [1], unmanned aerial vehicle (UAV)-assisted networking [2] and fifth-generation (5G) terrestrial and non-terrestrial network (NTN) services [3].
-All these applications require an accurate description of radio frequency (RF) spectrum coverage and efficient assessment of the radio environment, which lead to the new concept of radiomaps [4].
-Describing the RF power spectral density (PSD) over different geometric locations, frequencies, and time, radiomaps can enrich the information of spectrum coverage and assist the aforementioned spectrum management applications [5].
-Usually, dense radiomaps are reconstructed from sparse observations collected by mobile sensors and user devices, as shown in Fig.
-1.
-To fully leverage the power of radiomaps in spectrum resource allocation and network optimization, efficient radiomap estimation (RME) from sparse observations has emerged as an urgent challenge.
+Rapid growth of sensor networks and edge computing continue to advance the development of next-generation wireless communications, stimulating many novel technologies, such as vehicle-to-everything [1] and integrated sensing and communications [2].
+All these applications require an accurate description of radio frequency (RF) spectrum coverage and efficient assessment of the radio environment, leading to the concept of radiomaps [3].
+Radiomaps describe RF power spectral density (PSD) across various locations, frequencies, and time, offering rich information of spectrum coverage for the spectrum management applications.
+Usually, dense radiomaps are reconstructed from sparse observations collected by mobile sensors.
+To fully leverage the power of radiomaps in spectrum management, efficient radiomap estimation (RME) from sparse observations has emerged as an urgent challenge.
 Existing methods of radiomap estimation can be categorized into model-based or learning-based approaches, each bearing its distinct constraints.
-Model-based methods usually assume a specific radio propagation model, with which unknown parameters are estimated by minimizing the error between observed samples and predictions.
-Typical examples include the log-distance path loss (LDPL) models [6], inverse distance weighted interpolation [7] and thin-plate splines kernels [8].
-However, the performances of model-based RME highly rely on the selection of physical radio propagation models, which are usually unknown in practical applications.
+Model-based approaches, such as the log-distance path loss (LDPL) models [4], inverse distance weighted interpolation (IDW) [5], and thin-plate splines kernels [6], usually assume a specific radio propagation model, with which unknown parameters are estimated by minimizing the error between observed samples and predictions.
+However, in practical applications, these physics models are often unknown, making model selection challenging.
 Moreover, the PSD distribution sometimes is more sensitive to surrounding obstacles rather than following a strict empirical model, especially in complicated environments.
-The efficient extraction of radio model information from realistic data samples remains an open question.
+Thus, the efficient extraction of radio model information from realistic data samples remains an open question.
 With the development of artificial intelligence (AI), recent interests of RME focus on learning-based methods [7].
@@
 Another work utilizing deep convolutional neural networks is [9], where an auto-encoder structure is presented for radiomap estimation.
-Other learning-based methods also include the conditional generative adversarial nets (cGAN) [3] and transformers [12].
-Despite many successes, efficient learning-based RME requires sufficient data samples with satisfying quality, which are usually inaccessible in realistic scenarios.
-Although some works [13, 14] consider integrating the radio propagation models with learning machines to address the data scarcity, the efficient model selection remains challenging.
-Moreover, how to satisfy the special requirements of downstream tasks, such as outage detection and network planning, is another critical concern.
-To address the aforementioned problems, this work proposed a Task-incentivized Radiomap Estimation based on Generative Adversarial Networks (TiRE-GAN) for radiomap estimation, leveraging the synergies of radio propagation model and generative AI.
+Other learning-based methods also include conditional generative adversarial nets (cGAN) [10, 11] and transformers [12].
+However, these learning-based methods require high-quality, abundant data, which is usually inaccessible in practical applications.
+Although some hybrid methods [13, 14] attempt to combine radio models with machine learning approaches, the efficient model selection and task-specific designed RME, such as for outage detection and network planning, remain critical concerns.
+To address the aforementioned problems, we propose a Task-incentivized Radiomap Estimation using Generative Adversarial Networks (TiRE-GAN), leveraging the synergies of radio propagation model and generative AI.
 Specifically, a radio depth map is introduced as an additional input channel of the generator to capture the overall radio propagation information and shadowing effects.
@@
 Our contributions can be summarized as follows:
-  • We introduce a novel radio depth map (RDM) as input conditions to the cGAN, serving as a proxy for the shadowing effects and radio propagation behavior, thereby infusing the learning process with additional insights into the physical principles of radio propagation.
-  • We integrate the physics-conditioned cGAN with a novel task-incentivized network to provide compensation to the estimated radiomaps depending on downstream tasks, which significantly enhances the RME performance.
-  • Our experimental results demonstrate the power of the proposed radio depth map in characterizing radio propagation behavior, as well as the efficiency of the proposed TiRE-GAN in radiomap estimation.
+We introduce a novel radio depth map (RDM) as an input to the cGAN, capturing shadowing and radio propagation effects to enhance the learning process with physical insights.
+We develop a physics-conditioned cGAN with a task-incentivized network, which improves RME performance by compensating radiomaps based on downstream tasks.
+Experiments demonstrate the effectiveness of the proposed RDM in capturing radio propagation behavior and the efficiency of TiRE-GAN.
 II Problem Description
-Figure 1: Problem Description of RME: Given the distribution of buildings, transmitter locations (in read), and the sparse observations collected by mobile sensors and devices, we aim to predict the dense radiomaps.
-Figure 2: Overall Structure of TiRE-GAN consisting of 3 parts: 1) The model-based data preprocessing module utilizes the input observed information (e.i.
+Figure 1: Overall Structure of TiRE-GAN consisting of 4 parts: 1) The model-based data preprocessing module utilizes the input observed information (i.e.
 building map and transmitter location) to generate the radio depth map and then stacks the four layers together as input; 2) The radiomap (RM) generation module is in a cGAN generator structure, which learn and produce the prediction of radiomap; 3) The feedback module consists of 2 parts: a cGAN discriminator and a task-incentivized network.
-Both modules produce a loss term as feedback to improve the generator.
+Both modules produce a loss term as feedback to improve the generator; 4) The Downstream task that provides the additional loss term for training guidance.
 In this section, we first introduce our problem formulation.
-As shown in Fig.
-1, the radiomap estimation is formulated as a sparse sensing problem, where a dense radiomap $\hat{M}_{R}$ is reconstructed given the landscape information of the surrounding environment, sparse observations collected from sensors and mobile devices, and locations of transmitters, formulated by the following equation:
+In this work, the radiomap estimation is formulated as a sparse sensing problem, where a dense radiomap $\hat{M}_{R}$ is reconstructed given the landscape information of the surrounding environment, sparse observations collected from sensors, and locations of transmitters, formulated by the following equation:
 $\hat{M}_{R}=g(M_{U},M_{T},M_{S},\mathcal{X}),$ (1)
-where $M_{U}$, $M_{T}$, $M_{S}$ and $\mathcal{X}$ represent the landscape map, the encoded transmitter location, the sampled radiomap, and additional extracted features, respectively.
+where $g(\cdot)$ denotes the RME neural network.
+Here, $M_{U}$, $M_{T}$, and $M_{S}$ represent the landscape map, the encoded transmitter location, and the sampled radiomap, respectively.
+$\mathcal{X}=f_{physical}(M_{U},M_{T})$ is a function embedding the physical principles of radio propagation behavior.
+In this work, radio depth map is applied to capture the $f_{physcal}(\cdot)$, which will be introduced with details in Section III-B2.
 Different from the interpolation formulation in [4] and [5], where the unobserved PSD values are estimated only using the observed samples in the same radiomap, we follow the classic prediction formulation and assume a set of training pairs of $(M_{R},\mathcal{F})$ with =$\mathcal{F}=\{M_{U},M_{T},M_{S},\mathcal{X}\}$ to train the mapping function $g(\cdot)$ by
-$\min_{g(\cdot)}\quad L(M_{R},\hat{M}_{R}),$ (2)
+$\min_{g(\cdot)}\quad L(M_{R},\hat{M}_{R}),\vspace{-1mm}$ (2)
 where $L(\cdot)$ refers to the designed loss function.
-In this work, our objectives can be summarized as two parts: 1) design an effective mapping function $g(\cdot)$ by integrating the cGAN with a novel task-incentivized network; and 2) design a suitable loss function $L(\cdot)$ to embed the physical principles of radio propagation.
-We shall introduce the details of our proposed TiRE-GAN in Section III.
+In this work, our objectives are to: 1) design an effective mapping function $g(\cdot)$ by integrating the cGAN with a novel task-incentivized network; and 2) design a suitable loss function $L(\cdot)$ to embed the physical principles of radio propagation.
+We introduce the details of our proposed TiRE-GAN in Section III.
 III Methods
@@
 As shown in Fig.
-2, the architecture of the proposed framework can be structured into three components: 1) model-based data preprocessing; 2) radiomap generation based on cGAN; and 3) task-incentivized feedback.
-The overall operational sequence of the training process is delineated as follows.
-To embed the radio propagation model, the urban landscape map and the transmitter location data are processed to fabricate a physics-embedded Radio Depth Map (RDM), which will be further discussed in Section III-B2.
-Subsequently, this RDM, alongside the sampled radiomap, urban map, and transmitter location data, are amalgamated into a quadruple-channel input for the cGAN.
-Throughout the training phase, a pre-trained task-incentivized network is incorporated as a contributory factor within the loss function for the cGAN’s generator, enhancing the model’s ability to compensate for the radiomap estimations by embedding physics-based constraints and the rewards from downstream tasks.
-We will then introduce each module as follows.
+1, the architecture of the proposed framework can be structured into three components: 1) model-based data preprocessing; 2) radiomap generation; and 3) task-incentivized feedback.
+The training process operates as follows.
+First, the urban landscape map and transmitter location data are processed to construct a physics-embedded Radio Depth Map (RDM), detailed further in Section III-B2, after which it is fed into the cGAN generator together with the urban map, transmitter locations and sparse observations.
+During training, a pre-trained task-incentivized network is integrated to provide feedback by embedding physics-based constraints and rewards from downstream tasks.
+Each module will be introduced as follows.
 III-B Model-Based Data Preprocessing
@@
 Similar to the building map, we encode the transmitter locations into a binary image $M_{T}\in\mathbb{R}^{N\times N}$, where the transmitter locations are marked as 1; otherwise, the pixel values are zero.
-The sparse observations can be viewed as the corresponding pixels in gray-scale images $M_{S}\in\mathbb{R}^{N\times N}$.
+The sparse observations can be viewed as a grayscale image $M_{S}\in\mathbb{R}^{N\times N}$ via zero-padding, where we use the observed PSD the as corresponding pixel values; otherwise, the pixel value is zero.
 For convenience, we first use a threshold to eliminate the extremely low-value observations, and then normalize the PSD value into $0\sim 1$ for the observed locations.
@@
 For the missing values in $M_{S}$, we set them as zeros.
-Since a convolutional neural network is applied for the cGAN structure, zero-padding does not impact the final radiomap estimation.
+Since a convolutional neural network is applied for the cGAN structure, zero-padding does not impact the final RME.
 III-B2 Radio Depth Map
+In Eq.
+(1), we included an $\mathcal{X}$ of additional extracted features for RME.
 To embed the radio propagation information, we introduce the Radio Depth Map (RDM), $\mathcal{X}=M_{D}\in\mathbb{R}^{N\times N}$ as a novel model-based attribute, derived from the distribution of buildings and the spatial locations of transmitters.
 Similar to a depth map in computer vision which provides information on the distances of surfaces from a specific viewpoint, the RDM is formulated as:
-$M_{D}(x,y)=n\left(\sum_{t=1}^{N_{T}}P_{t}(x,y)\cdot B_{t}(x,y)\right),$ (3)
-where $(x,y)$ denotes the two-dimensional coordinates of the target location, $P_{t}(x,y)$ denotes the distance-based path loss gain from the $t$-th transmitter, and the $B_{t}(x,y)$ capture the impact of buildings from the viewpoint of $t$-th transmitter.
-Suppose that $C(l,l_{t})$ identifies the set of pixels in the direct line of sight between the target location $l=(x,y)$ and location $l_{t}=(x_{t},y_{t})$ of the $t$-th transmitter.
+$M_{D}(\xi,\eta)=n\left(\sum_{t=1}^{N_{T}}P_{t}(\xi,\eta)\cdot B_{t}(\xi,\eta)\right),$ (3)
+where $(\xi,\eta)$ denotes the two-dimensional coordinates of the target location, $N_{T}$ is the number of transmitters, $P_{t}(\xi,\eta)$ denotes the distance-based path loss gain from the $t$-th transmitter, $n(\cdot)$ is the max normalization function to re-scale the maximal value in the RDM as $1$, and the $B_{t}(\xi,\eta)$ capture the impact of buildings from the viewpoint of $t$-th transmitter.
+Let $C(\ell,\ell_{t})$ denote the set of pixels in the direct line of sight between the target location $\ell=(\xi,\eta)$ and location $\ell_{t}=(\xi_{t},\eta_{t})$ of the $t$-th transmitter.
 The shadowing term impacted by buildings is calculated by
-$B_{t}(x,y)=\frac{\sum_{i\in C(l,l_{t})}(1-M_{U}(i))}{\sum_{i\in C(l,l_{t})}1},$ (4)
-which characterizes the ratio of non-building pixels between $l$ and $l_{t}$. $B_{t}(x,y)$ is larger and the shadowing effect is smaller if the radio propagates through fewer buildings.
-To capture the path loss $P_{t}(x,y)$, we employ the Inverse Distance Weighting (IDW) model, calculated by
-$P_{t}(x,y)=d(l,l_{t})^{-\lambda},\ \lambda>0,$ (5)
-where $d(l,l_{t})$ denotes the Euclidean distance between the target point $l$ and the $t$-th the transmitter.
+$B_{t}(\xi,\eta)=\frac{\sum_{i\in C(\ell,\ell_{t})}(1-M_{U}(i))}{\sum_{i\in C(\ell,\ell_{t})}1},\vspace{-1mm}$ (4)
+which characterizes the ratio of non-building pixels between $\ell$ and $\ell_{t}$. $B_{t}(\xi,\eta)$ is larger and the shadowing effect is smaller if the radio propagates through fewer buildings.
+To capture the path loss $P_{t}(\xi,\eta)$, we employ the Inverse Distance Weighting (IDW) model, calculated by
+$P_{t}(\xi,\eta)=d(\ell,\ell_{t})^{-\lambda},\ \lambda>0,$ (5)
+where $d(\ell,\ell_{t})$ denotes the Euclidean distance between the target $\ell$ and the $t$-th transmitter.
 The parameter $\lambda$ governs the rate, at which signal strength diminishes with increasing distance, necessitating a positive value to ensure physical plausibility in the depth map.
 An exemplary radio depth map is shown in the orange block of Fig.
-2.
+1.
 III-C cGAN Network
-In this work, we employ a conditional Generative Adversarial Network (cGAN) [15] as the foundational architecture, which consists of two principal components: a generator $G$ and a discriminator $D$.
-Different from the traditional GAN which generates data from random noise, the generator in cGAN can generate data from given features $\mathcal{F}$ and incorporate additional information $\mathbf{I}$ in the discriminator.
-In our radiomap estimation, the conditional features consist of four channels as processed in the Data Preprocessing Module, i.e., =$\mathcal{F}=\{M_{U},M_{T},M_{S},M_{D}\}$ as shown in Fig.
-2.
-The additional information $\mathbf{I}$ can be designed as a function of the region features, i.e., $\mathbf{I}=h(\mathcal{F})$ or some determined labels, such as fake or real [16].
+In this work, we employ a cGAN [15] as the foundational architecture, which consists of two principal components: a generator $G$ and a discriminator $D$.
+The generator employs a ResNet-based architecture, consisting of 6 ResNet blocks.
+The discriminator uses a PatchGAN structure, comprising 4 convolutional blocks , and ends with a Sigmoid activation to classify the input as real or generated.
+In our radiomap estimation, the conditional features consist of four channels as processed in the model-based data preprocessing module, i.e., =$\mathcal{F}=\{M_{U},M_{T},M_{S},M_{D}\}$ as shown in Fig.
+1.
+The additional information $\mathbf{I}$ can be designed as a function of the region features, i.e., $\mathbf{I}=h(\mathcal{F})$ or some determined labels, such as surrounding landscape information [16].
 Then, the basic cGAN can be formulated as
@@
 $\mathbb{E}_{f\sim p_{\rm data}(\bf{\mathcal{F}})}[\log(1-D(G(\bf{\mathcal{F}}),\mathbf{I}))],$ (6)
-where $p_{\rm data}(\bf{y,I})$ is the joint distribution with the prior knowledge $\mathbf{I}$, and $\mathbf{y}$ is the generated data.
+where $p_{\rm data}(\bf{y,I})$ is the joint distribution with the prior knowledge $\mathbf{I}$, and $\mathbf{y}$ is the true radiomap.
+$p_{\rm data}(\mathbf{\mathcal{F}})$ denotes the distribution of input features used for cGAN training.
+Interested readers could refer to [16, 10] for a detailed introduction to cGAN structure.
 Generally, the optimization of a min-max problem in cGAN can be split into two sub-problems, and train $D$ and $G$ iteratively.
-Specifically, the loss function to optimize $D$ is
-$L_{D}=-V(D,G),$ (7)
-and the general loss function to optimize $G$ is
+Specifically, the loss function to optimize $D$ is $L_{D}=-V(D,G)$, and the general loss function to optimize $G$ is
 $L_{G}=\mathbb{E}_{x\sim p_{\rm data}(\bf{\mathcal{F}})}[\log(1-D(G(\bf{\mathcal{F}}),\mathbf{I}))].$ (7)
@@
 $L=L_{G}+\alpha L_{MSE}+\beta L_{R}.$ (8)
-Here, $L_{MSE}$ describes the error of RME by Mean Squared Error (MSE), i.e.,
-$L_{MSE}(\mathbf{y},\hat{\mathbf{y}})=\frac{1}{K}\sum_{i=1}^{K}(y_{i}-\hat{y}_{i})^{2},$ (10)
-where $K$ is the total number of pixels, $\mathbf{y}$ and $\hat{\mathbf{y}}$ are the ground truth and the predicted radiomap, while $y_{i}$ and $\hat{y}_{i}$ are the true and predicted pixel value of the $i$th pixel.
+Here, $L_{MSE}$ describes the error of RME by Mean Squared Error (MSE), i.e., $L_{MSE}(\mathbf{y},\hat{\mathbf{y}})=\frac{1}{K}\sum_{i=1}^{K}(y_{i}-\hat{y}_{i})^{2}$, where $K$ is the total number of pixels, $\mathbf{y}$ and $\hat{\mathbf{y}}$ are the ground truth and the predicted radiomap, while $y_{i}$ and $\hat{y}_{i}$ are the true and predicted pixel value of the $i$th pixel.
 $L_{R}$ is a task-incentivized regularization term from the feedback module, which will be introduced in Section III-D.
@@
 The discriminator is trained by the binary cross entropy (BCE) loss as
-$\vspace{-5mm}L_{D}(\mathbf{y},\hat{\mathbf{y}})=-\frac{1}{K}\sum_{i=1}^{K}\left[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right].$ (11)
-Figure 3: Task-incentivized network (TIN) structure: We design the TIN based on Unet, which contains 4 encoding blocks and 4 decoding blocks.
-Each encoding block consists of 3 convolution layers and a pooling layer.
+$L_{D}(\mathbf{\delta},\hat{\mathbf{\delta}})=-\frac{1}{K}\sum_{i=1}^{K}\left[\delta_{i}\log(\hat{\delta}_{i})+(1-\delta_{i})\log(1-\hat{\delta}_{i})\right],\vspace{-4mm}$ (9)
+where $\delta$ and $\hat{\delta}$ are the true and predicted values. $K$ is the number of data points in the batch.
+Figure 2: Task-incentivized network (TIN): based on Unet, TIN contains 4 encoding blocks and 4 decoding blocks.
+Each encoding block uses 3 convolution layers and a pooling layer.
 Each decoding block consists of one concatenation layer, a transpose convolution layer, and 3 convolution layers.
 III-D2 Task-Incentivized Network
-In traditional RME, the objective transcends the mere accuracy of radio signal strength at discrete locations within a dense radiomap.
-However, in practical scenarios, the final objective of RME is to assist spectrum planning and network management.
-For example, one can easily identify the outage of regions based on their radiomap by setting a certain threshold of service failure.
-To compensate for the radiomap estimations based on downstream tasks, we introduce a TIN module as shown in Fig.
+The goal of TiRE-GAN extends beyond capturing radio signal strength at specific locations; it shall ultimately support spectrum planning and network management.
+For instance, radiomaps can highlight service outages by setting a failure threshold.
+To enhance radiomap estimations for downstream tasks, we introduce a TIN module, as shown in Fig.
+1.
+We use outage detection as an exemplary TIN, employing a UNet architecture for semantic segmentation of radio power outages as shown in Fig.
 2.
-As an exemplary TIN, we adopt outage detection as a downstream task.
-More specifically, the proposed TIN utilizes a UNet architecture for semantic segmentation of radio power outage as Fig.
-3, where the input constitutes a radiomap and the output is an outage map indicating signal coverage or absence.
-The training of the TIN is conducted using ground-truth radio coverage and outage maps, thereby ensuring that the network learns to accurately predict outage areas based on real-world data.
-Once trained, the TIN operates within the cGAN framework as a pre-trained component.
-During inference, the output generated by the cGAN’s generator serves as the input to the TIN, enabling it to contribute an additional term $L_{R}$ to the loss function of the cGAN.
-This term is specifically designed to encapsulate aspects of human-centric interest, such as the distribution and characteristics of signal outages, thereby guiding the cGAN toward generating radiomaps that are not only precise but also aligned with the real-world considerations and requirements of downstream tasks.
-This integration of the TIN within the cGAN architecture exemplifies a strategic convergence of machine learning precision with domain-specific insights, enhancing the utility and applicability of the generated radiomaps.
-The TIN provides the task-incentivized regularization term $L_{R}$ in the loss function of the generator, which is calculated by the pre-trained TIN using the MSE loss to measure the error between the predicted and the ground truth outage maps:
-$L_{R}(z,\hat{z})=\frac{1}{K}\sum_{i=1}^{K}(z_{i}-\hat{z}_{i})^{2},$ (12)
-where the $z$ and $\hat{z}$ are the ground truth and the predicted outage map, while $z_{i}$ and $\hat{z}_{i}$ are the true and predicted pixel values of the $i$th pixel. $K$ is the number of pixels.
-Note that, our TIN can be integrated with any RM-assisted applications. Some other typical examples also include ray tracing or wireless localization.
-Figure 4: Performance comparison (sr=1) between Kriging, matrix completion(MC), AutoEncoder (AE), RadioUNet (RU), conditional GAN (cGAN), our TiRE-GAN, and the ground truth (GT) on 3 radiomap examples.
+The input is a radiomap, and the output outage map indicates signal coverage.
+Trained on real radio coverage and outage maps, the TIN accurately predicts outage areas.
+The TIN operates within the cGAN framework as a pre-trained component, for which the generator’s output serves as input.
+The TIN then contributes an additional loss term, $L_{R}$, guiding cGAN to produce radiomaps that not only maintain precision but also address real-world concerns relevant to downstream tasks.
+The TIN provides a task-incentivized regularization term $L_{R}$ in the loss function of the generator, which is calculated by the pre-trained TIN using the MSE loss to measure the error between the predicted and the ground truth outage maps, i.e.,
+$L_{R}(z,\hat{z})=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}(z_{i}-\hat{z}_{i})^{2},\vspace{-2mm}$ (10)
+where the $z$ and $\hat{z}$ are the ground truth and the predicted outage map, while $z_{i}$ and $\hat{z}_{i}$ are the true and predicted pixel values of the $i$th pixel.
+$N_{p}$ is the number of pixels.
+Note that our TIN can be integrated with any RM-assisted applications. Some typical examples include ray tracing or wireless localization.
+Figure 3: Performance comparison (sr=1) between Kriging, AutoEncoder (AE), RadioUNet (RU), conditional GAN (cGAN), RME-GAN, Geo2SigMap, our TiRE-GAN, and the ground truth (GT) on 3 radiomap examples.
 TiRE-GAN achieves the best radio propagation and the shadowing effect prediction.
@@
 IV-A Experiment Setup
-IV-A1 Data description
-In this work, we utilize the RadiomapSeer dataset [10] for evaluation, which comprises an extensive set of 700 maps from a variety of metropolitan areas, including Ankara, Berlin, Glasgow, Ljubljana, London, and Tel Aviv.
-Each map delineates a quadrant of $256\times 256$ square meters, depicted with a granularity of $256\times 256$ pixels.
+IV-A1 Data
+In this work, we utilize the RadiomapSeer dataset [8] for evaluation, which contain an extensive set of 700 maps from a variety of metropolitan areas, including Ankara, Berlin, Glasgow, Ljubljana, London, and Tel Aviv.
+Each map delineates a quadrant of $256\times 256$ square meters, at a granularity of $256\times 256$ pixels.
 The edifices are encoded by assigning a pixel intensity of one, whereas open spaces are represented by a zero value.
 Furthermore, the precise locations of transmitters are encoded with a pixel intensity of one superimposed on a null background, facilitating accurate localization.
-IV-A2 Method Comparison
-To evaluate the performance of our TiRE-GAN, we compare with both state-of-the-art (SOTA) model-based methods, including Kriging [17], and data-driven approaches, including Matrix completion (MC) [18], AutoEncoder (AE) [11], RadioUnet (RU) [10] and conventional cGAN [15].
-For the interpolation-based method, we reconstruct the radiomap with observations in the same radiomap.
-For the prediction-based method, we split the dataset as 4000/800/800 (70%/15%/15%) for training/validation/testing sets.
+IV-A2 Comparison
+For performance assessment, we compare TiRE-GAN with both state-of-the-art (SOTA) model-based methods, including Kriging [17], and data-driven approaches, including AutoEncoder (AE) [9], RadioUnet (RU) [8], conventional cGAN [15], RME-GAN [10], and Geo2SigMap [18].
+For interpolation-based method, we reconstruct the radiomap with observations in the same radiomap.
+For prediction-based method, we split the dataset as 4000/800/800 (70%/15%/15%) for training/validation/testing sets.
+Since RME-GAN is designed for non-uniformly distributed samples, we remove the corresponding loss terms to ensure fair comparison.
 In our TiRE-GAN, we set $\alpha=1$ and $\beta=0.1$ in Eq.
@@
 IV-B Overall performance
-We first present the overall performance of different radiomap estimations under different sampling ratios (sr).
-In this work, we evaluate the performance based on the normalized radiomaps via MSE denoted by
-$\text{MSE}(\mathbf{y},\hat{\mathbf{y}})=\frac{1}{K}\sum_{i=1}^{K}(y_{i}-\hat{y}_{i})^{2},$ (13)
-and normalized MSE (NMSE) calculated by
-$\text{NMSE}(\mathbf{y},\hat{\mathbf{y}})=\frac{1}{K}\sum_{i=1}^{K}\frac{(y_{i}-\hat{y}_{i})^{2}}{y_{i}^{2}},$ (14)
-where $K=256^{2}$.
-Note that, since the linear normalization is applied, we can easily re-scale the estimated radiomap back to the original data domain.
-The results are presented in Table I and II.
-From the results, our proposed TiRE-GAN always achieves superior performance against all other SOTA approaches.
-Specifically, with the assistance of our proposed radio depth map, our method can achieve a very high reconstruction accuracy even with a very small sampling ratio and insufficient training samples, which validates the effectiveness of the proposed method.
-The visualization results of different methods in 3 radiomap examples are shown in Fig.
-4, where the sample ratio is 1%.
-From the visualization results, the radioUnet, cGAN, and TiRE-GAN have better performances than Kringing, MC, and AE in capturing the overall radio propagation patterns.
-Compared to radioUnet and cGAN, our TiRE-GAN has more explicit boundaries for the shadowing effects, resulting from the compensation from the TIN feedback.
-The numerical results, together with the visualization results demonstrate the efficiency of our proposed TiRE-GAN for radiomap estimation.
-TABLE I: Test NMSE Comparison (×$\times 10^{-2}$)
-sr (%) 0.03 0.1 0.5 1 3 5 10 Average
-Krigin 24.5696 18.9280 14.2451 12.9608 14.7171 11.3540 10.9647 15.3913
-RU 6.5421 5.1111 2.6374 2.2206 1.8192 1.5191 1.2704 2.8744
-AE 54.7656 53.8440 45.6640 43.3302 26.0041 20.7167 14.3275 36.9503
-AE deep 50.6597 48.8197 21.4272 16.2537 6.8252 6.5503 5.8179 156.3537
-cGAN 11.5754 5.4805 2.1345 1.8301 1.3009 1.1894 1.0295 3.5057
-MC 83.0969 74.7810 69.0133 66.9996 62.8215 56.3645 55.4314 66.9297
-TiRE-GAN 1.9866 1.2608 0.7282 0.6084 0.5228 0.5119 0.4615 0.8686
-TABLE II: Test MSE Comparison (×$\times 10^{-4}$)
-sr (%) 0.03 0.1 0.5 1 3 5 10 Average
-Krigin 188.7554 141.3767 110.7954 101.5476 91.9940 89.0443 86.0208 115.5340
-RU 24.6458 19.3893 10.0221 8.4693 7.0856 5.9640 4.9978 11.5105
-AE 230.6564 226.6141 194.4702 186.5724 113.0439 87.7205 65.0134 157.7272
-AE deep 209.2621 201.4013 91.9589 70.0012 26.8177 26.4321 23.4719 92.7636
-cGAN 45.8797 22.0908 8.4572 7.2467 5.0700 4.6442 4.0541 13.9203
-MC 342.1892 307.9448 284.1937 275.9010 258.6960 232.1065 228.2640 275.6136
-TiRE-GAN 7.0604 4.6453 2.7678 2.3472 2.0377 2.0015 1.8137 3.2390
+We first present the overall performance of different radiomap estimations.
+We evaluate the performance based on the normalized radiomaps via MSE denoted by $\text{MSE}(\mathbf{y},\hat{\mathbf{y}})=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}(y_{i}-\hat{y}_{i})^{2}$, and normalized MSE (NMSE) calculated by $\text{NMSE}(\mathbf{y},\hat{\mathbf{y}})=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}\frac{(y_{i}-\hat{y}_{i})^{2}}{y_{i}^{2}}$, where $N_{p}=256^{2}$.
+Linear normalization makes it easy to re-scale the estimated radiomap back to the original data domain.
+From the results given in Table I, our proposed TiRE-GAN always achieves superior performance against all other approaches.
+Different from RME-GAN and purely data-driven models, our radio depth map can capture both pathloss and shadowing efficiently, providing efficient guidance of cGAN training.
+With the assistance of radio depth map, our method can achieve a very high reconstruction accuracy even for very low sampling ratio and few training samples.
+Fig.
+3 presents the visualization results of different methods for three radiomap examples at a 1% sample ratio.
+Advanced learning methods, such as RadioUnet, cGAN, and our TiRE-GAN, outperform Kringing and AE in capturing overall radio propagation patterns.
+Notably, due to the model-based interpolation block and complex loss design, RME-GAN cannot efficiently function at very low sampling ratio below 0.1%.
+TiRE-GAN shows clearer boundaries for shadowing effects resulted from the TIN feedback.
+Both the visualization and numerical results validate the effectiveness of our TiRE-GAN in RME with insufficient training samples.
+TABLE I: Test NMSE (×$\times 10^{-2}$)/MSE (×$\times 10^{-4}$) Comparison
+sr (%) 0.03 0.1 0.5 1 3 5 10 Average ($0.5\sim 10$)
+Kriging 24.56/188.75 18.92/141.37 14.24/110.79 12.96/101.54 14.71/91.99 11.35/89.04 10.96/86.02 12.84/95.88
+AE 54.76/230.65 53.84/226.61 45.66/194.47 43.33/186.57 26.00/113.04 20.71/87.72 14.32/65.01 30.00/129.36
+AE Deep 50.65/209.26 48.81/201.40 21.42/91.95 16.25/70.00 6.82/26.81 6.55/26.43 5.81/23.47 11.22/47.73
+RU 6.54/24.64 5.11/19.38 2.63/10.02 2.22/8.46 1.81/7.08 1.51/5.96 1.27/4.99 1.89/7.30
+cGAN 11.57/45.88 5.48/22.09 2.13/8.46 1.83/7.25 1.30/5.07 1.18/4.64 1.02/4.05 1.49/5.89
+RME-GAN -/- -/- 0.77/7.42 0.68/5.70 0.64/6.15 0.53/5.08 0.47/4.51 0.60/5.77
+Geo2SigMap 2.19/8.67 1.23/4.73 1.12/4.31 0.75/2.89 0.79/2.93 0.79/2.94 0.95/3.75 0.93/3.36
+TiRE-GAN 1.98/7.06 1.26/4.65 0.72/2.77 0.61/2.35 0.52/2.04 0.51/2.00 0.46/1.81 0.55/2.19
 IV-C Robustness Against Noise
-To evaluate the robustness of TiRE-GAN, we conduct the experiment, where a noisy input pattern is substituted for the original sparse observation channel by introducing additive Gaussian noise to each sampled value in the radiomap.
-To showcase the effectiveness of the proposed radio depth map and TIN blocks, we compare our TiRE-GAN with Kriging, matrix completion(MC), AutoEncoder (AE), RadioUNet (RU), and conditional GAN (cGAN) under different signal-to-noise ratios (SNR) as Fig.
-5.
-Here, we consider a sampling ratio of 1% for all the test cases.
-From the results, our TiRE-GAN always has better performance.
-Moreover, the MSE of TiRE-GAN increases slower than those of other methods, especially compared to the high performance methods (RadioUNet, cGAN), demonstrating the robustness of the TiRE-GAN against noise.
-Figure 5: The MSE (×$\times 10^{-4}$) of Kriging, matrix completion(MC), AutoEncoder (AE), RadioUNet (RU), conditional GAN (cGAN), and TiRE-GAN of different SNRs under a sample ratio=1%.
+To assess the robustness, we include additive Gaussian noise to each sampled radiomap value, replacing the original sparse observation channel with a noisy input pattern.
+To highlight the effectiveness of the proposed radio depth map and TIN blocks, we compared different methods across various signal-to-noise ratios (SNR), as shown in Fig.
+4, using a 1% sampling ratio for all test cases.
+TiRE-GAN consistently outperforms the other methods.
+Moreover, its MSE growth is slow, particularly compared to advanced learning methods like RadioUNet and cGAN, which further demonstrates robustness of TiRE-GAN against noise.
+Figure 4: The MSE (×$\times 10^{-4}$) comparison with different SNRs.
 IV-D Outage Evaluation of Generated radiomap
-To further evaluate the performance of reconstructed radiomaps in downstream tasks, we compare the performance of RM-assisted outage detection.
+To further evaluate the effect on downstream tasks, we compare the performance of RM-assisted outage detection.
 Here, we input the generated radiomaps from different methods into the pre-trained outage detection networks.
-The numerical results of the whole datasets with sr=1% and the exemplary visualization results are shown in Table III and Fig.
-6, respectively.
-From the results, the traditional model-based methods and conventional auto-encoder perform poorly in the outage detection.
-Our TiRE-GAN always has superior performance, which demonstrates the effectiveness of TIN in radiomap compensation for downstream tasks.
-TABLE III: Outage Map MSE comparison
-Method Kriging MC AE
-MSE 0.078134 0.075765 0.049377
-Method RU cGAN TiRE-GAN
-MSE 0.028439 0.030352 0.026682
+The numerical results of the whole datasets with 1% sampling and the exemplary visualization results are shown in Table II and Fig.
+5, respectively.
+RME-GAN and TiRE-GAN achieve lower MSE than other methods, resulted from their utilization of model information.
+However, in the visualization, TiRE-GAN displays smoother and more consistent patterns, which should benefit the spectrum management.
+These results demonstrate the effectiveness of TIN in radiomap compensation.
+TABLE II: Outage Map MSE comparison
+Method Kriging AE AEdeep RU
+MSE 0.0781 0.0758 0.0494 0.0284
+Method cGAN Geo2SigMap RME-GAN TiRE-GAN
+MSE 0.0304 0.0561 0.0231 0.0267
 Figure 5: Illustration of Outage Map Prediction under sr=1%.
-We compared the outage map predicted by the generated radiomap from Kriging, matrix completion(MC), AutoEncoder (AE), RadioUNet (RU), conditional GAN (cGAN), and our TiRE-GAN.
+We compared the outage map predicted by the generated radiomap from different methods.
 The corresponding radiomap and the ground truth outage map(GT) are shown on the right for reference.
-The result shows that TiRE-GAN has the best capability to predict the outage map with a clear boundary and range.
 V Conclusion
-In this work, we introduce a task-incentivized generative learning model, namely TiRE-GAN for radiomap estimation.
-More specifically, to characterize the overall radio propagation behavior, we introduce a radio depth map as additional physics-embedded features as the condition to guide the data-driven generative models.
-To compensate for the radiomap estimation for downstream tasks, we introduce a task-incentivized network to provide feedback to predict the detailed features in the radiomap.
-From our experimental results, the proposed TiRE-GAN always achieves superior performance, even with insufficient training samples, which addresses the barriers of conventional learning approaches.
-Our numerical and visualization results demonstrate the effectiveness of the proposed radio depth map and TiRE-GAN.
-Future research endeavors will delve into the optimization of the generator’s reward mechanism.
-Another promising direction is the multi-band radiomap estimation for different frequencies.
+TiRE-GAN is a task-incentivized generative learning model for radiomap estimation.
+To capture the overall radio propagation behavior, it uses a radio depth map as a physics-embedded feature to guide the data-driven generative model.
+Additionally, a task-incentivized network provides feedback to enhance detailed radiomap features for downstream tasks.
+Experimental results show that TiRE-GAN consistently outperforms other methods, especially with limited training samples, a faced by realistic applications.
+Future work will focus on optimizing the generator’s reward mechanism and exploring multi-band radiomap estimation across different frequencies.